OpenStudy (anonymous):

A candle burns 4 inches in 100 minutes. What is the unit rate

OpenStudy (dean.shyy):

Inches/Minute?

OpenStudy (anonymous):

Would it be \[\frac{ 4 }{ 100 }=\frac{ 2 }{ 50 } \frac{ inches }{ \min }\] ?

OpenStudy (anonymous):